# Mosaic IT: Enhancing Instruction Tuning with Data Mosaics

### v1

@@ Abstract
Finetuning large language models with a variety of instruction-response pairs has enhanced their capability to understand and follow instructions. Current instruction tuning primarily relies on teacher models or human intervention to generate and refine the instructions and responses, which are costly, non-sustainable, and may lack diversity. In this paper, we introduce Mosaic Instruction Tuning (Mosaic-IT), a human/model-free method that can efficiently create rich and diverse augmentations from existing instruction tuning data to enhance the finetuned LLM.Mosaic-IT randomly concatenates multiple instruction data into one and trains the model to produce the corresponding responses with predefined higher-level meta-instructions to strengthen its multi-step instruction-following and format-following skills. Our extensive evaluations demonstrate a superior performance and training efficiency of Mosaic-IT, which achieves consistent performance improvements over various benchmarks and a $80\%$ reduction in training costs compared with original instruction tuning. Our codes and data are available at https://github.com/tianyi-lab/Mosaic-IT.
@@ 1 Introduction
The emergence of Large Language Models (LLMs) [1, 41, 39, 45, 46, 19] along with their remarkable performance in down-stream tasks [64, 55, 47, 57, 32, 34, 56, 63, 31, 4, 43, 59], has revolutionized the domains of Artificial Intelligence and Natural Language Processing. A key component of the recipe to unlock the exceptional ability of LLMs in understanding and following instructions is the technique of Instruction Tuning (IT) [38, 52, 8, 51, 62, 55], which involves the fine-tuning of LLMs on datasets comprising corresponding instruction-response pairs.
@@
To ensure the quality of instruction tuning data, earlier efforts [1, 39, 45, 19] carefully curate extensive, diverse, and high-quality datasets manually. Although these datasets encompass a wide range of instructions to improve instruction tuning, they require the responses to be meticulously curated by human experts [20, 58, 52, 50, 14]. Alternatively, some approaches [49, 44, 54, 25, 55] leverage more capable teacher LLMs to reduce the labor-intensive process of data generation. For example, the Alpaca [44] utilizes self-instruct [49] to automatically generate diverse instruction tuning datasets, and the WizardLM [54] proposes to complicate the existing instruction data by an evolution algorithm. Building on this trend and the widely acknowledged notion that more complicated instructions are more beneficial for LLMs’ instruction-following ability [54, 65], numerous strategies [65, 53, 12, 25, 33, 24, 23, 17, 55] have been proposed to further diversify and complexify the instruction-response pairs, utilizing teacher models like ChatGPT-3.5 and GPT-4 [39].
@@
Despite the enhanced performance in instruction-following ability offered by these existing methods, they face Two major limitations: (1) They heavily rely on teacher models to rewrite instruction-response pairs, which highlights the resource-intensive nature and their constraints on scalability; (2) They only increase the complexity within the scope of a single instruction, which limits the potential improvement in LLMs’ instruction-following capabilities. Rethinking the current methodologies, we hypothesize that the process of instruction tuning should not be constrained by one single instruction but be extended to follow several instructions at a time, which represents a higher level of instruction-following ability that is beneficial to the training process. A similar concept during the inference phase is proposed by batch prompting [5], where multiple samples are grouped in one batch allowing LLMs to generate multiple responses at one inference. However, its sub-optimal performances further indicate the complexity of this setting and the necessity of further training for this higher-level capability.
@@
As orthogonal to the existing instruction tuning methods, we introduce Mosaic Instruction Tuning (Mosaic-IT), an innovative and model-free approach that augments existing instruction tuning datasets, which concurrently improves the LLM performances and lowers the training expenses. Specifically, in our method, multiple instructions and corresponding responses from the original dataset are concatenated into a single sample for fine-tuning, simulating the multi-instruction-following scenarios at no cost. Without applying any additional strategies, we term this simple process as the Primary Mosaic Strategy. Our empirical findings illustrate that this simple data augmentation, devoid of any supervision costs, proves remarkably effective. We posit that this mosaic strategy process significantly improves the complexity of the original instructions, learning from which directly benefits LLMs in their instruction-following ability. Additionally, this method offers the advantage of directly reducing the total count of instruction-response pairs, thereby cutting down on training iterations, and accelerating the training process significantly by approximately $80\%$ reduction.
@@
Though effective, the Primary Mosaic strategy constrains LLMs in responding to the instructions in the original order and format, potentially limiting its further potential. Thus we further introduce three Auxiliary Mosaic Strategies aimed at enhancing the diversity and complexity of the mosaicked instruction-response pairs: Format, Permute, and Maskout, in which an additional meta-instruction is provided as a higher-level guideline for LLMs to follow the given instructions. Specifically, in the Format strategy, some arbitrary parsing formats will be defined in the meta-instruction thus forcing LLMs to follow these formats, which notably enhances the LLMs’ capacity to follow formats. In the Permutation strategy, an arbitrary permuted order is defined thus forcing LLMs to respond in a desired order. In the Maskout strategy, some arbitrary instructions are sampled which meta-instruction forces LLMs to ignore. Moreover, the use of these auxiliary strategies not only boosts the performance in several evaluation metrics but also keeps our method free of additional costs.
@@
  • We extend existing instruction tuning on one instruction at a time into following multiple instructions in diverse forms, which dramatically digs out the potential of the existing datasets.
  • We introduce an innovative Mosaic-IT approach for instruction tuning, supported by comprehensive evaluations showcasing its superior performance.
  • Notably, the Mosaic process is free of cost, eliminating the necessity and cost of collecting additional supervision from humans or teacher models. Moreover, our Mosaic-IT approach significantly enhances training efficiency by reducing required training iterations.
2 Related Work
Earlier research in instruction tuning primarily centered on constructing expansive, high-quality datasets through intensive curation by human experts, a process both time-consuming and labor-intensive [20, 58, 52, 50, 14]. Motivated by the success of Alpaca [44], recent studies have explored automated approaches for developing instruction-tuning datasets.
Instruction Data Improvement: WizardLM [54] first proposes an Evol Algorithm to complicate the existing data and reach supreme performance. LaMini-LM [53] innovatively generates "Topic-Guided" instructions utilizing Wiki data. Tree-Instruct [65] preliminarily explores the relationship between instruction complexity and Alignment and proposes adding nodes to complicate the instruction. UltraChat [12] establishes broad thematic scopes, systematically generating numerous instructions within each. Reflection-Tuning [25] sequentially refines both instructions and responses by focusing on specific evaluative criteria. DEITA [33] utilizes ChatGPT to diversify and then select the data. Selective Reflection-Tuning [24] proposes a teacher-student collaborative pipeline to improve and select the data. Instruction Fusion [17] proposes to utilize ChatGPT4 to merge two distinct instructions for further complexity enhancement. These advancements showcase a shift towards automating the generation and refinement of datasets, reducing reliance on human labor.
Instruction Data Selection: It is widely accepted that "quality is all you need" [46, 67] for instruction tuning. LIMA [67] demonstrates that merely 1,000 human-carefully-curated, high-quality training instances can substantially enhance the instruction-following performance. InsTag [37] employs the proprietary model, ChatGPT, to tag instruction data and select data with complex tags. Alpagasus [3] utilizes proprietary LLMs chatGPT and Claude2 to directly assess the quality of instruction tuning data. Cherry LLM [27] proposes the Instruction-Following Difficulty (IFD) scores to assess the difficulty of the instructions, which is a self-guided method in which no extra LLMs are utilized. Motivated by Humpback [28], Selective Reflection-Tuning [24] extends the IFD score to a reverse version, focusing on the feasibility of responses. [13] and [2] utilize reward models as the base scores for measuring data quality. DEITA [33] experiments on several different data selection metrics and builds a dataset with high quality. Superfiltering [26] reveals the consistency between weak and strong language models in perceiving instruction difficulty, making the filtering process much more efficient. All these works are devoted to distinguishing and selecting good data samples from bad ones for instruction tuning.
@@ 3 Methodology
Figure 1: Illustration of Mosaic-IT. The upper section represents the original Instruction Tuning method. In the middle is our method with the Primary Mosaic strategy, which concatenates instructions and corresponding responses as an augmentation for better Instruction Tuning. On the right are corresponding data examples of each method. The lower section illustrates the three Auxiliary Mosaic Strategies for Mosaic-IT.
@@ 3.1 Preliminaries
The instruction tuning dataset, defined as $D$, consists of $n$ data samples, each represented by a triplet $(\textit{Instruction},[\textit{Input}],\textit{Response})$. Historically, instruction tuning data has emphasized separate segments for both the instruction and input components to enhance control [50, 36, 44]. In contrast, contemporary datasets often integrate the input directly into the instruction for simplicity [67, 7, 54, 25, 24]. Following this trend, we define $x=\texttt{map}(\textit{Instruction},[\textit{Input}])$ as the unified instruction, and $y$ as the corresponding response. The mapping function can be as simple as concatenation, incorporating specific control tokens. Therefore, $D$ can be represented as ${(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{n},y_{n})}$, denoting a set of $n$ instruction-response pairs. Let $p_{\theta}(\cdot)$ denote the LLMs to be trained, with parameters $\theta$. In the instruction tuning setting, $p_{\theta}$ is typically fine-tuned by maximizing the following objective on each data $(x_{i},y_{i})$, $y_{i,j}$ represents the $j_{th}$ token of response $y_{i}$ and $l_{i}$ represents the token length:
$\max_{\theta}\sum_{j=1}^{l_{i}}\log p_{\theta}\left(y_{i,j}|x_{i}\right),$ (1)
@@ 3.2 Mosaic-IT
Motivated by the success of the existing data-centric instruction tuning methods, a line of approaches is proposed to further enhance the instruction-response pairs utilizing extra teacher LLMs. Though effective, all existing methods for instruction tuning restrict training samples to just one instruction, which severely limits the potential of the existing high-quality data and the instruction-following ability of the models to be trained. We hypothesize that the process of instruction tuning should not be constrained by one single instruction but be extended to follow several instructions at a time, which represents a higher level of instruction-following ability that is beneficial to the training process. Specifically, we propose Mosaic Instruction Tuning (Mosaic-IT) as shown in Figure 1, where the middle section illustrates our method with the Primary Mosaic strategy, and the lower section illustrates the three Auxiliary Mosaic Strategies.
@@ 3.2.1 Primary Mosaic strategy
Exploring the concept of concatenating unrelated instruction-response pairs into a unified instruction-response pair for training remains largely unexplored. The primary challenge lies in crafting a coherent overall instruction and acquiring the corresponding response. Most existing methods utilize a strong teacher model to rewrite and polish the instructions with prompting techniques and generate corresponding responses, introducing more cost by actually re-generating new data samples. To harness the full potential of existing data rather than directly replacing them, we introduce a simple stacking approach as shown in Figure 1 (middle), in which instructions are randomly concatenated alongside serial digits to form an overall instruction. The concatenated overall instruction is denoted as $[x_{1},...,x_{k}]$, with the corresponding overall response concatenated as $[y_{1},...,y_{k}]$. Here, $k$ denotes the number of original data samples integrated into each overall sample.
@@
In this framework, the fundamental instruction-following capability is grounded in the existing instruction-response pairs, and the mosaic strategy extends this capability to a higher level in which LLMs are forced to follow multiple instructions. It represents a much more complicated scenario that benefits LLMs compared with traditional single-task instructions. Consequently, the objective function for each concatenated overall data sample can be formulated as follows:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y_{1},...,y_{k}]_{j}|[x_{1},...,x_{k}]\right),$ (2)
@@
Here, $[y_{1},...,y_{k}]_{j}$ denotes the $j{th}$ token of the overall response, and $l$ represents its length. This formulation encapsulates the essence of our approach, optimizing the model parameters $\theta$ to maximize the likelihood of generating the correct sequence of responses for the given overall instruction.
3.2.2 Auxiliary Mosaic Strategies
@@
Though effective, this simple Primary Mosaic strategy constrains LLMs in responding to the instructions with the original order and format, potentially limiting its generalization and practical usage. In our method, the instructions and corresponding responses from the original dataset can be viewed as atomic components and our method randomly combines these elements together to form new instructions and responses. This nature allows us to further complicate this process with fancier strategies thus forcing LLMs to follow more complicated overall instructions. Hence, we propose three Auxiliary Mosaic Strategies to complicate and diversify the mosaicked samples as shown in Figure 1 (lower), including Format, Permute, and Maskout, with meta-instructions guiding them.
@@
Format In the Format strategy, some arbitrary formats are defined in the meta-instruction to force LLMs to follow these formats. The formats mainly contain two categories: 1) Serial Digit Format and (2) Response Parsing Format. The serial digits establish the initial instruction order that guides LLMs to follow sequentially. We manually define $10$ types of serial digit format, which will be randomly sampled during each mosaic process. For response parsing, we simulate the scenario where the users try to extract specific information from the responses. We define $27$ types of parsing brackets and $17$ types of parsing text pairs, which will be randomly sampled and assembled during each mosaic process. Examples can be found in Appendix C, which can be easily extended for customized training settings. We denote responses with specific formats as $y^{\prime}_{i}=wrap(y_{i},s_{format})$, and $l$ as the token length of the overall response. An additional meta-instruction $s_{format}$ specifying the required format will be included in the overall instruction. Thus, the objective function for each mosaic data point:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y^{\prime}_{1},...,y^{\prime}_{k}]_{j}|[x_{1},...,x_{k},s_{format}]\right),$ (3)
@@
In the Permute strategy, an arbitrary permuted order is defined in the meta-instructions, forcing LLMs to follow. Moreover, several high-level rules are defined to ensure the complexity and diversity of meta-instructions, e.g., forcing LLMs to respond to each instruction in the randomly generated permutation list, forcing LLMs to respond in the alphabetical order of each instruction, forcing LLMs to respond according to the length of instructions, etc. The detailed rule types and descriptions are depicted in Appendix C. These various meta-instructions not only provide higher-level guidelines for LLMs to follow multiple instructions but also inherently enhance the instruction perception ability of LLMs. The meta-instruction is denoted as $s_{permute}$ and is included in the overall instruction. The permuted response list is denoted as $[y^{\prime}_{1^{\prime}},...,y^{\prime}_{k^{\prime}}]=Permute([y^{\prime}_{1},...,y^{\prime}_{k}],s_{permute})$. Thus the objective function can be formulated as below:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y^{\prime}_{1^{\prime}},...,y^{\prime}_{k^{\prime}}]_{j}|[x_{1},...,x_{k},s_{format},s_{permute}]\right),$ (4)
In the Maskout strategy, some arbitrary instructions are selected in the meta-instructions forcing LLMs to ignore them. Several high-level rules are also defined similarly to the permute strategy, including forcing LLMs to ignore the instructions with given random digits, forcing LLMs to ignore the longest one/several instructions, forcing LLMs to ignore odd-numbered instructions, etc. The details are provided in Appendix C. Similarly, the meta-instruction is denoted as $s_{maskout}$ and the response list is denoted as $[y^{\prime}_{1},...,y^{\prime}_{m}]=Maskout([y^{\prime}_{1},...,y^{\prime}_{k}],s_{maskout})$, where $m$ is the count of responses after masking out. Thus the objective function can be formulated as below:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y^{\prime}_{1},...,y^{\prime}_{m}]_{j}|[x_{1},...,x_{k},s_{format},s_{maskout}]\right)$ (5)
It’s important to note that our mosaic strategies entail no supervision cost, and the predefined rules are flexible and have the potential for further extension. We utilize the version with three auxiliary strategies as our default Mosaic-IT.

### v2

# Mosaic-IT: Free Compositional Data Augmentation Improves Instruction Tuning

@@ Abstract
Finetuning large language models with a variety of instruction-response pairs has enhanced their capability to understand and follow instructions. Current instruction tuning primarily relies on teacher models or human intervention to generate and refine the instructions and responses for training, which are costly, non-sustainable, and may lack diversity. In this paper, we introduce Mosaic Instruction Tuning (Mosaic-IT), a human/model-free compositional data augmentation method that can efficiently create rich and diverse augmentations from existing instruction tuning data to enhance the LLMs. Mosaic-IT randomly concatenates multiple instruction data into one and trains the model to produce the corresponding responses with predefined higher-level meta-instructions to strengthen its multi-step instruction-following and format-following skills. Our extensive evaluations demonstrate a superior performance and training efficiency of Mosaic-IT, which achieves consistent performance improvements over various benchmarks and a $80\%$ reduction in training costs compared with original instruction tuning. Our codes and data are available at https://github.com/tianyi-lab/Mosaic-IT.
@@ 1 Introduction
The emergence of Large Language Models (LLMs) Brown et al. (2020); Scao et al. (2022); OpenAI (2023); Touvron et al. (2023a; b); Jiang et al. (2023) along with their remarkable performance in down-stream tasks Zhao et al. (2023); Xu et al. (2024a), has revolutionized the domains of Artificial Intelligence and Natural Language Processing. A key component of the recipe to unlock the exceptional ability of LLMs in understanding and following instructions is the technique of Instruction Tuning (IT) Mishra et al. (2021); Wei et al. (2022); Chung et al. (2022); Wang et al. (2023c); Zhang et al. (2023); Xu et al. (2024a), which involves the fine-tuning of LLMs on datasets comprising corresponding instruction-response pairs.
@@
To ensure the quality of instruction tuning data, earlier efforts Brown et al. (2020); OpenAI (2023); Touvron et al. (2023a); Jiang et al. (2023) carefully curate extensive, diverse, and high-quality datasets manually. Although these datasets encompass a wide range of instructions to improve instruction tuning, they require the responses to be meticulously curated by human experts Khashabi et al. (2020); Ye et al. (2021); Wei et al. (2022); Wang et al. (2022); Du et al. (2022). Alternatively, some approaches Wang et al. (2023b); Taori et al. (2023); Xu et al. (2023); Li et al. (2023a) leverage more capable teacher LLMs to reduce the labor-intensive process of data generation. For example, the Alpaca Taori et al. (2023) utilizes self-instruct Wang et al. (2023b) to automatically generate diverse instruction tuning datasets, and the WizardLM Xu et al. (2023) proposes to complicate the existing instruction data by an evolution algorithm. Building on this trend and the widely acknowledged notion that more complicated instructions are more beneficial for LLMs’ instruction-following ability Xu et al. (2023); Zhao et al. (2024), numerous strategies Zhao et al. (2024); Wu et al. (2024); Ding et al. (2023); Li et al. (2023a); Liu et al. (2023a); Li et al. (2024b; a); Guo et al. (2024); Xu et al. (2024a) have been proposed to further diversify and complexify the instruction-response pairs, utilizing teacher models like ChatGPT-3.5 and GPT-4 OpenAI (2023).
@@
Despite the enhanced performance in instruction-following ability offered by these existing methods, they face Two major issues: (1) They heavily rely on teacher models or human annotators to rewrite instruction-response pairs, which highlights the resource-intensive nature and their constraints on scalability; (2) They only increase the complexity within the scope of a single instruction, which limits the potential improvement in LLMs’ instruction-following capabilities. Motivated by the Dense and Aligned Captions Doveh et al. (2023) proposed for vision language (VL) models and the mosaic data augmentation proposed in Yolov4 Bochkovskiy et al. (2020), we hypothesize that denser instructions benefit the LLM alignment, i.e. the process of instruction tuning should not be constrained by one single instruction but be extended to follow several instructions at a time, which represents a higher level of instruction-following ability that is beneficial to the training process. A similar concept during the inference phase is proposed by batch prompting Cheng et al. (2023); Lin et al. (2024), where multiple samples are grouped in one batch allowing LLMs to generate multiple responses at one inference, while its performances are sub-optimal. Moreover, our preliminary experiments on GPT-3.5-turbo and GPT-4-turbo show that even for these strong proprietary LLMs, their performances degrade dramatically if required to follow several instructions at one time, the experimental results are presented in the Section 5 Further Discussion. Thus, these performance degradation phenomenons indicate the complexity of this setting and the necessity of further training for this higher-level capability.
Figure 1: The illustration of our Mosaic-IT with different strategies. Given the original dataset, our method randomly samples and concatenates them together into more complex samples, simulating the multi-instruction-following scenarios at no cost.
@@
As orthogonal to the existing instruction tuning methods, we introduce Mosaic Instruction Tuning (Mosaic-IT), an innovative and model/human-free compositional approach that augments existing instruction tuning datasets, which concurrently improves the LLM performances and lowers the training expenses. As shown in Figure 1, in our method, multiple instructions and corresponding responses from the original dataset are concatenated into a single sample for fine-tuning, simulating the multi-instruction-following scenarios at no cost. Without applying any additional strategies, we term this simple process as the Primary Mosaic Strategy. We posit that this mosaic strategy process significantly improves the complexity and density of the original instructions, learning from which directly benefits LLMs in their instruction-following ability. Additionally, this method offers the advantage of directly reducing the total count of instruction-response pairs, thereby cutting down on training iterations, and accelerating the training process significantly by approximately $80\%$ reduction.
@@
Though effective, the Primary Mosaic strategy constrains LLMs in responding to the instructions in the original order and format, potentially limiting its further potential. Thus we further introduce three Advanced Mosaic Strategies aimed at enhancing the diversity and complexity of the mosaicked instruction-response pairs: Format, Permute, and Maskout, in which an additional meta-instruction is provided as a higher-level guideline for LLMs to follow the given instructions. Illustrative examples are presented in Figure 2. Specifically, in the Format strategy, some arbitrary parsing formats will be defined in the meta-instruction thus forcing LLMs to follow these formats, which notably enhances the LLMs’ capacity to follow formats. In the Permutation strategy, an arbitrary permuted order is defined thus forcing LLMs to respond in a desired order. In the Maskout strategy, some arbitrary instructions are sampled which meta-instruction forces LLMs to ignore. Moreover, the use of these Advanced strategies not only boosts the performance in several evaluation metrics but also keeps our method free of additional costs.
@@
  • We propose a novel human/model-free data augmentation method, Mosaic-IT, which extends existing instruction tuning from handling one single instruction at a time to following multiple instructions in diverse forms. This approach significantly enhances the potential utilization of existing high-quality datasets.
  • Mosaic-IT improves the instruction-following abilities of LLMs compared to training on original data, as evidenced by consistent performance gains across a wide range of benchmarks, model families, and datasets, demonstrating strong generalization capabilities.
  • Mosaic-IT substantially increases training efficiency by reducing the required number of training iterations, resulting in an approximate 80% reduction in training time, as confirmed by experimental results.
Figure 2: Illustrative examples of Mosaic-IT. Given $3$ simple data points, our method can concatenate them into overall data samples with diverse forms. Texts in red represent the meta-instructions that define the formats or orders for LLMs to respond. Texts in yellow are major response differences of each strategy. The Primary Strategy only concatenates data together. The Format Strategy requires LLMs to respond in predefined formats. The Permute Strategy requires LLMs to respond in specific orders and the Maskout Strategy requires LLMs to ignore some of the instructions.
@@ 2.1 Preliminaries
[…]
For simplicity, we define $x=\texttt{map}(\textit{Instruction},\textit{Input})$ as the unified instruction, and $y$ as the corresponding response. Therefore, $D$ can be represented as ${(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{n},y_{n})}$, denoting a set of $n$ instruction-response pairs. Let $p_{\theta}(\cdot)$ denote the LLMs to be trained, with parameters $\theta$. In the instruction tuning setting, $p_{\theta}$ is typically fine-tuned by maximizing the following objective on each data $(x_{i},y_{i})$, $y_{i,j}$ represents the $j_{th}$ token of response $y_{i}$, $y_{i,<j}$ represents the tokens prior to $y_{i,j}$, and $l_{i}$ represents the token length of $y_{i,j}$:
$\max_{\theta}\sum_{j=1}^{l_{i}}\log p_{\theta}\left(y_{i,j}|x_{i},y_{i,<j}\right),$ (1)
@@ 2.2 Mosaic-IT
Motivated by the success of the existing data-centric instruction tuning methods, a line of approaches is proposed to further enhance the instruction-response pairs utilizing extra teacher LLMs Xu et al. (2024a). Though effective, all existing methods for instruction tuning restrict training samples to just one instruction, which severely limits the potential of the existing high-quality data and the instruction-following ability of the models to be trained. Motivated by the Dense and Aligned Captions Doveh et al. (2023) for VL, we hypothesize that denser instructions benefit the LLM alignment, thus the process of instruction tuning should not be constrained by one single instruction but be extended to follow several instructions at a time, which represents a higher level of instruction-following ability that is beneficial to the training process. Thus, we propose Mosaic Instruction Tuning (Mosaic-IT) as shown in Figure 1.
@@ 2.2.1 Primary Mosaic strategy
Exploring the concept of concatenating random instruction-response pairs into a unified instruction-response pair for training remains largely unexplored. The primary challenge lies in crafting a coherent overall instruction and obtaining its corresponding response. Most existing methods utilize a strong teacher model to rewrite and polish the instructions with prompting techniques and generate corresponding responses, introducing more cost by actually re-generating new data samples. To harness the full potential of existing data rather than directly discarding them, we introduce a simple compositional approach as shown in Figure 2, in which instructions are randomly concatenated with serial digits to form an overall instruction. The concatenated overall instruction is denoted as $[x_{1},...,x_{k}]$, with the corresponding overall response concatenated as $[y_{1},...,y_{k}]$. Here, $k$ denotes the number of original data samples integrated into each overall sample.
@@
In this framework, the fundamental instruction-following capability is triggered by the existing instruction-response pairs, and the mosaic strategy extends this capability to a higher level in which LLMs are forced to follow multiple instructions. It represents a much more complicated scenario that benefits LLMs compared with traditional single-task instructions. Consequently, the objective function for each concatenated overall data sample can be formulated as follows:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y_{1},...,y_{k}]_{j}|[x_{1},...,x_{k}],[y_{1},...,y_{k}]_{<j}\right),$ (2)
@@
Here, $[y_{1},...,y_{k}]_{j}$ denotes the $j{th}$ token of the overall response, $[y_{1},...,y_{k}]_{<j}$ denotes the tokens prior to $j{th}$ token, and $l$ represents the length of overall response. This formulation encapsulates the essence of our approach, optimizing the model parameters $\theta$ to maximize the likelihood of generating the correct sequence of responses for the given overall instruction.
2.2.2 Advanced Mosaic Strategies
@@
Though effective, this simple primary mosaic strategy constrains LLMs in responding to the instructions with the original order and format, potentially limiting its generalization and practical usage. In our method, the instructions and corresponding responses from the original dataset can be viewed as atomic components and our method randomly combines these elements together to form new instructions and responses. This nature allows us to further complicate this process with fancier strategies thus forcing LLMs to follow more complicated overall instructions. Hence, we propose three Advanced Mosaic Strategies to complicate and diversify the mosaicked samples as shown in Figure 2, including Format, Permute, and Maskout, with meta-instructions guiding them.
@@
Format In the Format strategy, some arbitrary formats are defined in the meta-instruction to force LLMs to follow these formats in the response. The formats mainly contain two categories: 1) Serial Digit Format and (2) Response Parsing Format. The serial digits establish the initial instruction order that guides LLMs to follow sequentially. We manually define $10$ types of serial digit format, which will be randomly sampled during each mosaic process. For response parsing, we simulate the scenario where the users try to extract specific information from the responses. We define $27$ types of parsing brackets and $17$ types of parsing text pairs, which will be randomly sampled and assembled during each mosaic process. Examples can be found in Appendix D, which can be easily extended for customized training settings. We denote responses with specific formats as $y^{\prime}_{i}=wrap(y_{i},s_{format})$, and $l$ as the token length of the overall response. An additional meta-instruction $s_{format}$ specifying the required format will be included in the overall instruction. Thus, the objective function for each mosaic data point:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y^{\prime}_{1},...,y^{\prime}_{k}]_{j}|[x_{1},...,x_{k},s_{format}],[y^{\prime}_{1},...,y^{\prime}_{k}]_{<j}\right)$ (3)
@@
In the Permute strategy, an arbitrary permuted order is defined in the meta-instructions, forcing LLMs to follow. Moreover, several high-level rules are defined to ensure the complexity and diversity of meta-instructions, e.g., forcing LLMs to respond to each instruction in the randomly generated permutation list, forcing LLMs to respond in the alphabetical order of each instruction, forcing LLMs to respond according to the length of instructions, etc. The detailed rule types and descriptions are depicted in Appendix D. These various meta-instructions not only provide higher-level guidelines for LLMs to follow multiple instructions but also inherently enhance the instruction perception ability of LLMs. In our settings, LLMs are required to generate responses selectively conditioned on some critical parts of the overall instruction, forcing them to first understand the formats and other requirements, indicating a more comprehensive understanding of the context given. The meta-instruction is denoted as $s_{permute}$ and is included in the overall instruction. The permuted response list is denoted as $[y^{\prime}_{1^{\prime}},...,y^{\prime}_{k^{\prime}}]=Permute([y^{\prime}_{1},...,y^{\prime}_{k}],s_{permute})$. Thus the objective function can be formulated as below:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y^{\prime}_{1^{\prime}},...,y^{\prime}_{k^{\prime}}]_{j}|[x_{1},...,x_{k},s_{format},s_{permute},[y^{\prime}_{1^{\prime}},...,y^{\prime}_{k^{\prime}}]_{<j}]\right),$ (4)
In the Maskout strategy, some arbitrary instructions are selected in the meta-instructions forcing LLMs to ignore them. Several high-level rules are also defined similarly to the permute strategy, including forcing LLMs to ignore the instructions with given random digits, forcing LLMs to ignore the longest one/several instructions, forcing LLMs to ignore odd-numbered instructions, etc. The details are provided in Appendix D. Similarly, the meta-instruction is denoted as $s_{maskout}$ and the response list is denoted as $[y^{\prime}_{1},...,y^{\prime}_{m}]=Maskout([y^{\prime}_{1},...,y^{\prime}_{k}],s_{maskout})$, where $m$ is the count of responses after masking out. Thus the objective function can be formulated as below:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y^{\prime}_{1},...,y^{\prime}_{m}]_{j}|[x_{1},...,x_{k},s_{format},s_{maskout}],[y^{\prime}_{1},...,y^{\prime}_{m}]_{<j}\right)$ (5)
It’s important to note that our mosaic strategies entail no supervision cost, and the predefined rules are flexible and have the potential for further extension. We utilize the version with three Advanced strategies as our default Mosaic-IT.